OpenStudy (anonymous):

1.Which exponent makes the statement true? 1/5^9=5^? a)9 b)-9 c)1/9 d)-1/9 2.y^5 * y^14= a)2y^19 b)y^60 c)y^19 d)2y^60

OpenStudy (anonymous):

im not

OpenStudy (anonymous):

can some one help me

OpenStudy (anonymous):

stoppp

OpenStudy (anonymous):

Plug in the numbers that are in the answers

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

@e.cociuba do you know?

OpenStudy (aayushi.somani):

1) answer is b.. nd in 2) answer is a..!!

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

1) x^-n = 1/x^n

OpenStudy (aayushi.somani):

ya i m shure..!!

OpenStudy (anonymous):

i dont get it though

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

2) y^m * y^n = y^m+n

OpenStudy (aayushi.somani):

be more specific that what actually u didnt understand in these ques..??

OpenStudy (anonymous):

Do the math !

OpenStudy (chmvijay):